OpenStudy (anonymous):

solve each system by substitution: -x-y-3z=-9, z=-3x-1, x=5y-z+23

OpenStudy (amorfide):

well your second equation gives you the value of z substitute z into the other two equations now you have two equations to solve so you would then make y the subject using one of the two equations you have then substitute y into the other equation then you will get a value of x then substitute your value of x into the equation to get y then substitute x and y into original equations to get z

OpenStudy (amorfide):

let me start you off -x-y-3z=-9 equation 1 z=-3x-1 equation 2 x=5y-z+23 equation 3 sub equation 2 into 1 and 3 -x-y-3(-3x-1)=-9 equation 1 x=5y-(-3x-1)+23 equation 3 now simplify and follow my instructions from before